OpenStudy (anonymous):

x^2 cos^2 y - sin y = 0 find the derivative, I'm seriously having trouble with this problem, any help would be appreciated.

OpenStudy (zarkon):

\[2x\cos^2( y)+x^22\cos(y)(-\sin(y))y' - \cos(y)y' = 0\] solve for y'

OpenStudy (anonymous):

how did you get there? this is where I am at \[x ^{2}*-2cosy siny+2xcos ^{2}y -\cos y=0\]

OpenStudy (zarkon):

I assume you are doing implicit differentiation ... therefore every time you differentiate y with respect to x you get y' or \[\frac{dy}{dx}\]

OpenStudy (anonymous):

I am, just having trouble understand how you got from A to B. Your syntax is just confusing me.

OpenStudy (anonymous):

understanding*

OpenStudy (zarkon):

you have the same thing as me except you are missing the y'

OpenStudy (anonymous):

okay, well this is basically my question then, how do i differentiate cos^2 y and sin y

OpenStudy (zarkon):

treat every y as y(x) (it is a function of x) so when you take the derivative you need to use the chain rule

OpenStudy (zarkon):

\[\frac{d}{dx}\sin(y)=\cos(y)\frac{dy}{dx}\]

OpenStudy (anonymous):

so \[d/dx(\cos^2 y) \neq -2\cos y \sin y\]

OpenStudy (zarkon):

\[\frac{d}{dx}(\cos^2(x))=-2\cos(x)\sin(x)\] but \[\frac{d}{dx}(\cos^2(y))=-2\cos(y)\sin(y)\frac{dy}{dx}\]

OpenStudy (anonymous):

\[y'=\frac{2 x \cos (y)}{2 x^2 \sin (y)+1} \]

OpenStudy (anonymous):

could you just show me the steps of getting from cos^2 y to - 2 cos y sin y dy/dx

OpenStudy (zarkon):

bring down the 2 ...keep the inside the same (the cosine) ...drop the power down by 1...multiply by the derivative of the inside ...derivative of cosine is -sine...keep the inside of that the same (which is y)...then again by the chain rule multiply by the derivative of the inside...the derivative of y is dy/dx

OpenStudy (anonymous):

bring down the 2?

OpenStudy (zarkon):

\[\frac{d}{dx}(f(x))^2=2(f(x))^1f'(x)\] bring the two from the exponent ... out to the front.

OpenStudy (anonymous):

The attachment show how I solved it.

OpenStudy (anonymous):

...?

OpenStudy (zarkon):

you could also let \[F(x,y)=x^2 \cos^2 y - \sin y \] and compute \[-\frac{\frac{\partial F}{\partial x}}{\frac{\partial F}{\partial y}}\] ;)

OpenStudy (anonymous):

okay I guess my misunderstanding is how to apply the chain rule to a function with a y with in it. I have been working on the same problem for the past hour, this is in no way efficient learning. (Cos y)^2 lets say that y = u^2 u = cos m and m = y lets also say that y prime = 2 u u prime = -sin m and m prime = dy/dx the chain rule states that the derivative of f(g(x)) = fprime( g(x)) * gprime of (x) so 2 (cos m) * - sin (m) thats just the chain rule now how do I apply the last m ? Your explanations are a little vague for me to understand, I learn by example, if you were to just break it down to a walkthrough that would be fantastic, btw thanks for your time.

OpenStudy (zarkon):

\[\cos^2(y)=\cos^2(y(x))\] \[\frac{d}{dx}\cos^2(y(x))=2(\cos(y(x))^1(-\sin(y(x))y'(x))\] \[=-2\cos(y)\sin(y)y'\]

OpenStudy (anonymous):

why is there a superscript 1?

OpenStudy (anonymous):

I don't understand why this is confusing me so much. :(

OpenStudy (anonymous):

well, I give up.... I suppose I'll give you a medal for your effort.... :(